The perimeter of a rectangle is 50 feet. Describe the possible lengths of a side if the area of the rectangle is greater than 144 square feet.
Answer by mananth(16946)

The perimeter of a rectangle is 50 feet. Describe the possible lengths of a side if the area of the rectangle is greater than 144 square feet.
..
L*W>144..........1
2*(L+W)=50
L+W=25
L=25-w
...
plug value of L in equation 1
(25-w)*W>144
25W-W^2>144
W^2-25W+144>0
W^2-16W-9W+144>o
W(W-16)-9(W-16)>0
(W-16)(W-9)>0
When width is 9 then length should be>16
When width is 16 length >9
...
m.ananth@hotmail.ca